Show your solution and provide units.

1. Calculate the volume (in mL) of 2.25 M HNO3 solution required to react with 24.2 mL of 3.50 M NaOH solution.

2.What volume (in mL) of 4.35 M HCl solution is needed to react with 18.5 g of CA(OH)2 ?

3.How many grams of silver chromate (Ag2CrO4) will precipitate when 86.0 mL of 0.200 M silver nitrate (AgNO3) is added to 50.0 mL of 0.300 M chromate (K2CrO4)? ( Molar mass of Ag2CrO4 = 331.74 g/ mol).


1
Expert's answer
2021-10-15T02:55:46-0400

1. HNO3 + NaOH → NaNO3 + H2O

Proportion:

3.50 mol – 1000 mL

x mol – 24.2 mL

x = n(NaOH) "= \frac{3.50 \times 24.2}{1000} = 0.0847 \;mol"

According to the reaction equation:

n(HNO3) = n(NaOH) = 0.0847 mol

Proportion:

4.35 mol – 1000 mL

0.0847 – y mL

V(HNO3) "= \frac{0.0847 \times 1000}{4.35} = 19.47\;mL"

2. 2HCl + Ca(OH)2 → CaCl2 + 2H2O

M(Ca(OH)2) = 74.09 g/mol

n(Ca(OH)2) "= \frac{18.5}{74.09} = 0.249 \;mol"

According to the reaction equation:

n(HCl) = 2n(Ca(OH)2) "= 2 \times 0.249 = 0.498 \;mol"

Proportion:

4.35 mol – 1000 mL

0.498 mol – x mL

x = V(HCl) "= \frac{0.498 \times 1000}{4.35} = 114.5 \;mL"

3. 2AgNO3 + K2CrO4 → Ag2CrO4 + 2KNO3

n(AgNO3) "= \frac{0.200 \times 86.0}{1000} = 0.0172 \;mol"

n(K2CrO4) "= \frac{0.300 \times 50.0}{1000} = 0.015 \;mol"

Accordinf to the reaction for one mole of K2CrO4 we need two moles of AgNO3.

So, AgNO3 is the limiting reactant.

n(Ag2CrO4) = "\frac{1}{2}" n(AgNO3) "= \frac{1}{2} \times 0.0172 = 0.0086 \;mol"

m(Ag2CrO4) "= 0.0086 \times 331.74 = 2.85 \;mol"